Jupiter rotates once about its axis in 9 h 56 min. Its radius is 7.13 × 104 km. Imagine that you could somehow stand on the surface (although in reality that would not be possible, because Jupiter has no solid surface).a) Calculate your radial acceleration in meters per second squared. b)Calculate your radial acceleration in Earth g's.
Answer #1

given,
radius, R = 7.13*10^4 km

= 7.13*10^7 m

Time period, T = 9 hours 56 min

= 9*60*60 + 56*60

= 35760 s

we know, angular velocity , w = 2*pi/T

= 2*pi/35760

= 1.757*10^-4 rad/s

a) radial acceleration, a_rad = R*w^2

= 7.13*10^7*(1.757*10^-4)^2

= 2.2 m/s^2

b) a_rad = 2.2*g/9.8

= 0.2245*g